Show that simple harmonic motion is isochronous
1
Expert's answer
2019-12-09T08:46:17-0500

For simple harmonic motion

"\frac{d^2x}{dt^2}+\omega^2x=0"

"x=A\cos{(\omega t+\phi)}"

We can see that


"\omega=const"

Thus, simple harmonic motion is isochronous.